# Homeomorphism theorem for sums of translates on the real axis

### v1

@@ 3. Behavior of sums of translates for large arguments
In this section, we prove a key lemma which allows us to reduce our problem on the axis to the case of the segment. Another form of this statement was proven for sums of translates with positive multiples of a single kernel in [6, Lemma 4.3]. Now we are dealing with several kernels and we need a slightly modified estimate, so we will provide a proof.
@@
Suppose that a function $g$ is concave, nondecreasing on the semiaxis $M,\infty)$ and is continuous at $M.$ Then $g$ is uniformly continuous on $M,\infty)$.
@@
Applying Lemma 2.2, we obtain $\lim\limits_{\mathbf{x}\to\mathbf{y},\ t\to\infty}F(\mathbf{x},t)=-\infty.$
@@
$\forall M\in\mathbb{N},\ M>N\quad\exists\,\mathbf{y}_{M}\in\overline{S^{N}}\quad\exists\,t_{M}\in\mathbb{R}\quad\left(t_{M}<-M\quad\&\quad F(\mathbf{y}_{M},t_{M})>m_{0}(\mathbf{y}_{M})-1\right).$
Take $x\in(-\infty,-N)$. By our assumption, we have for all $k$ with $M_{k}<x<-N$
$J(x)=F(\mathbf{y}_{M_{k}},x)-\sum\limits_{j=1}^{n}K_{j}(x-y^{M_{k}}_{j})\leqslant F(\mathbf{y}_{M_{k}},t_{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(x-y^{M_{k}}_{j}).$
Using continuity of $K$ at $x-y^{*}_{j}<0$ and Lemma 3.1, we get
$J(x)$ $\leqslant\lim\limits_{k\to\infty}\left(F(\mathbf{y}_{M_{k}},t_{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(x-y^{M_{k}}_{j})\right)$
$=\lim\limits_{k\to\infty}F(\mathbf{y}_{M_{k}},t_{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(x-y^{*}_{j})=-\infty.$
So, $J(x)\equiv$-$\infty$ for x<N.x<-N. We have a contradiction with our assumption, since then also $F(\mathbf{y}_{M},t_{M})=-\infty.$
@@ 4. Local homeomorphism
    The functions $m_{j}:\overline{S^{\infty}}\to\underline{\mathbb{R}},\ j=0,\ldots,n,$ are continuous in the extended sense.
@@
    Without loss of generality, we can assume that the statement of Lemma 3.2 is true for $\tau_{N}-1$. Then for $\mathbf{y}\in\overline{S^{N}}\cap R^{N}$ we get
@@ 5. Properness of the difference function
The idea of the proof of the following lemma was communicated to us by Szilárd Gy. Révész. We already used this idea in [6, Th. 3.1] with his permission. However, we now prove a statement of a different nature, and therefore we need a slightly different estimate in the proof. Moreover, we are dealing with several kernels satisfying ($GM$) instead of just one monotone kernel.
@@
1. Denote $\mathbf{y}^{N}:=(y^{N}_{1},\ldots,y^{N}_{n}),\ y^{N}_{0}:=-\infty,\ y^{N}_{n+1}:=\infty.$ If $y^{N}_{n}\to\infty,$ take $i:=\min\{j:y^{N}_{j}\to\infty\}.$ Otherwise, we have $y^{N}_{1}\to$-$\infty$ and take $i:=\max\{j:y^{N}_{j}\to-\infty\}.$ Without loss of generality, consider the first case. The proof for the second case is similar.
@@
By definition of $i$, we have either $y^{N}_{i-1}\to$-$\infty$ or $y^{N}_{i-1}\to y_{i-1}\in\mathbb{R}.$ In both cases, there exist $A>0,\ N_{0}\in\mathbb{N}$ such that $y^{N}_{i-1}<A-1$ for all $N>N_{0}.$ Since $\{\mathbf{y}^{N}\}\subset R$ for all $N$, there are arbitrarily large points where $J$ is finite. Fix one such point $z>A.$ Then, for all $N>N_{0}$, we have $y^{N}_{i-1}<z-1$. Moreover, there exists $N_{1}>N_{0}$ such that $y^{N}_{i}>z+1$ for all $N>N_{1}$. For $N>N_{1},$ we conclude that $z\in(y^{N}_{i-1},y^{N}_{i})$ and $|z-y^{N}_{j}|>1$ for all $j.$ Next we consider these $N$.
@@ 3. Let us estimate
By Corollary 4.2, the functions $m_{j}$ are continuous. Therefore, the differences $m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})$ also satisfy $|m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})|\leqslant C$ for all $j=1,\ldots,n$. Taking into account the finiteness of $m_{i}(\mathbf{y})$, this implies that $\mathbf{y}\in R.$
@@ 6. Proof of the main result
is a local homeomorphism by Corollary 4.2 and it is proper by Lemma 5.2. Therefore, by Theorem C, we obtain that $D$ is a global homeomorphism. ∎
@@ 7. Homeomorphism theorem for sums of translates
Note that $K_{j},\ j=1,\ldots,n,$ and $J$ satisfy the conditions of Theorem 1.1. Therefore, the difference function $D:R\to\mathbb{R}^{n}$ for $F$ is a homeomorphism, and it is locally bi-Lipschitz. It remains to note that
@@ 8. Homeomorphism theorem for weighted generalized
Let $r_{1},\ldots,r_{n}>0$ be arbitrary. Denote $\mathbf{r}:=(r_{1},\ldots,r_{n})$. Consider the following set of monic generalized nonnegative polynomials [21, p. 392] of degree $r\leavevmode\nobreak\ :=\leavevmode\nobreak\ \sum\limits_{j=1}^{n}r_{j}$
@@
Let $K_{j}:=r_{j}\log|\cdot|,\ j=1,\ldots,n,\ J:=\log w.$ Obviously, these kernels and field satisfy the conditions of Theorem 1.1. Therefore, the difference function
@@
$\mathbf{y}\mapsto\left(\log\dfrac{\sup\limits_{t\in(y_{j},y_{j+1})}w(t)p(\mathbf{y},t)}{\sup\limits_{t\in(y_{j-1},y_{j})}w(t)p(\mathbf{y},t)}\right)_{j=1}^{n}$
@@ 9.1. Abstract log-concave interpolation on the real axis
Let
$K_{j}(t)$ $:=\log L_{j}(t),\quad j=1,\ldots,n,$
$J(t)$ $:=\begin{cases}\log w(x_{j}),\quad t=x_{j},\ j=0,\ldots,n,\\ -\infty,\quad t\in\mathbb{R}\setminus\{x_{0},\ldots,x_{n}\}.\end{cases}$
@@
By construction of $J$, the regularity set has the following form
$R:=\{\mathbf{y}\in\overline{S^{\infty}}:\ x_{j}<y_{j+1}<x_{j+1}\text{ for }j=0,\ldots,n-1\}.$
@@ 9.2. Moving node Hermite–Fejér interpolation
For convenience, let $y_{0}:=$-$\infty$ and $y_{n+1}:=\infty$. For any $\alpha_{0},\ldots,\alpha_{n}>0$ there are a unique $C>0$ and points $y_{1}<\ldots<y_{n}$ such that for the function
@@
The further proof is similar to the proof of Theorem 9.1 with $x_{j}$ replaced by $z_{j},\\ j=1,\ldots,n$. ∎

### v2

@@ 3. Behavior of sums of translates for large arguments
In this section, we prove a key lemma which allows us to reduce our problem on the axis to the case of the segment. Another form of this statement was proven for sums of translates with positive multiples of a single kernel in [6, Lemma 4.3]. Now we are dealing with several kernels and we need a slightly modified estimation for kernels, so we will provide a proof.
@@
Suppose that a function $g$ is concave, non-decreasing on the semiaxis $M,\infty)$ and is continuous at $M.$ Then $g$ is uniformly continuous on $M,\infty)$.
@@
Applying Lemma 2.2, together with the fact that for all $j\in J_{2},\ \lim\limits_{t\to\infty}K_{j}(t-c)$ is finite or -$\infty$, i.e., it is strictly less that $\infty,$ we obtain $\lim\limits_{\mathbf{x}\to\mathbf{y},\ t\to\infty}F(\mathbf{x},t)=-\infty.$
@@
$\forall M\in\mathbb{N},\ M>N\quad\exists\,\mathbf{y}^{M}\in\overline{S^{N}}\quad\exists\,t^{M}\in\mathbb{R}\quad\left(t^{M}<-M\quad\&\quad F(\mathbf{y}^{M},t^{M})>m_{0}(\mathbf{y}^{M})-1\right).$
We have the bounded sequence $\{\mathbf{y}^{M}\}\subset\overline{S^{N}}.$ By the Bolzano-Weierstrass theorem, there is a convergent subsequence $\{\mathbf{y}^{M_{k}}\}.$ Set
$\lim\limits_{k\to\infty}\mathbf{y}^{M_{k}}=:\mathbf{y}^{*}=(y^{*}_{1},\ldots,y^{*}_{n})\in\overline{S^{N}}.$
Take $t\in(-\infty,-N)$. By our assumption, we have for all $k$ with $M_{k}<t<-N$
$J(t)=F(\mathbf{y}^{M_{k}},t)-\sum\limits_{j=1}^{n}K_{j}(t-y^{M_{k}}_{j})\leqslant F(\mathbf{y}^{M_{k}},t^{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(t-y^{M_{k}}_{j}).$
Using continuity of $K_{j}$ at $t-y^{*}_{j}<0$ and Lemma 3.1, we get
$J(t)$ $\leqslant\lim\limits_{k\to\infty}\left(F(\mathbf{y}^{M_{k}},t^{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(t-y^{M_{k}}_{j})\right)$
$=\lim\limits_{k\to\infty}F(\mathbf{y}^{M_{k}},t^{M_{k}})+1-\sum\limits_{j=1}^{n}K_{j}(t-y^{*}_{j})=-\infty.$
So, $J(t)\equiv$-$\infty$ for t<N.t<-N. We have a contradiction with our assumption, since then also $F(\mathbf{y}^{M},t^{M})=-\infty.$
@@ 4. Local homeomorphism
    The functions $m_{j}:\overline{S^{\infty}}\to\underline{\mathbb{R}},\ j=0,\ldots,n,$ are continuous in the extended sense, i.e., as functions with values on the extended real axis.
@@
    Without loss of generality, we can assume that the statement of Lemma 3.2 is true for $\tau_{N}-1$. Then for $\mathbf{y}\in\overline{S^{N}}\cap R^{N}$ and $m_{i}^{\tau_{N}}(\mathbf{y})=m_{i}(\mathbf{y}),$ for $i=0,\ldots,n,$ we get
@@ 5. Properness of the difference function
The idea of the proof of the following lemma was communicated to us by Szilárd Gy. Révész. We already used this idea in [6, Th. 3.1] with his permission. However, we now prove a statement of a different setting, and therefore we need a slightly different estimate in the proof. Moreover, we are dealing with several kernels satisfying ($GM$) instead of just one monotone kernel.
@@
1. Denote $\mathbf{y}^{N}:=(y^{N}_{1},\ldots,y^{N}_{n}),\ y^{N}_{0}:=-\infty,\ y^{N}_{n+1}:=\infty.$ Since $\mathbf{y}^{N}$ is convergent in the extended sense, then $y^{N}_{n}\to\infty$, or $y^{N}_{1}\to$-$\infty$. Without loss of generality, we may assume the first case occurs, since the proof of the second case is similar. Let $i:=\min\{j:y^{N}_{j}\to\infty\}.$
@@
By definition of $i$, we have either $y^{N}_{i-1}\to$-$\infty$ or $y^{N}_{i-1}\to y_{i-1}\in\mathbb{R}.$ In both cases, there exist $A>0,\ N_{0}\in\mathbb{N}$ such that $y^{N}_{i-1}<A-1$ for all $N>N_{0}.$ Since $\{\mathbf{y}^{N}\}\subset R$ for all $N$, there are arbitrarily large points where $J$ is finite. Fix one such point $z>A.$ Then, for all $N>N_{0}$, we have $y^{N}_{i-1}<z-1$. Moreover, since $y^{N}_{i}\to\infty$ as $N\to\infty$, there exists $N_{1}>N_{0}$ such that $y^{N}_{i}>z+1$ for all $N>N_{1}$. For $N>N_{1},$ we conclude that $z\in(y^{N}_{i-1},y^{N}_{i})$ and $|z-y^{N}_{j}|>1$ for all $j.$ Next we consider these $N$.
@@ 3. Let us estimate
By Lemma 4.2, the functions $m_{j}$ are continuous. Therefore, the differences $m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})$ also satisfy $|m_{j}(\mathbf{y})-m_{j-1}(\mathbf{y})|\leqslant C$ for all $j=1,\ldots,n$. Taking into account the finiteness of $m_{i}(\mathbf{y})$, this implies that $\mathbf{y}\in R.$
@@ 6. Proof of the main result
is a local homeomorphism by Lemma 4.2 and it is proper by Lemma 5.2. Therefore, by Theorem C, we obtain that $D$ is a global homeomorphism. ∎
@@ 7. Homeomorphism theorem for sums of translates
Note that $K_{j},\ j=1,\ldots,n,$ and $J$ satisfy the conditions of Theorem 1.1. Therefore, the difference function $D:R\to\mathbb{R}^{n}$ for $F$ is a homeomorphism, and it is locally bi-Lipschitz. Moreover, the definition of $J$ implies that
@@ 8. Homeomorphism theorem for weighted generalized
Let $r_{1},\ldots,r_{n}>0$ be arbitrary. Denote $\mathbf{r}:=(r_{1},\ldots,r_{n})$. Consider the following set of monic generalized nonnegative polynomials [21, p. 392] of degree $r\penalty 10000\ :=\penalty 10000\ \sum\limits_{j=1}^{n}r_{j}$
@@
Let $K_{j}:=r_{j}\log|\cdot|,\ j=1,\ldots,n,\ J:=\log w.$ Obviously, these kernels and this field satisfy the conditions of Theorem 1.1. Therefore, the difference function
@@
$\mathbf{y}\mapsto\left(\log\dfrac{\sup\limits_{t\in(y_{j},y_{j+1})}w(t)p(\mathbf{y},t)}{\sup\limits_{t\in(y_{j-1},y_{j})}w(t)p(\mathbf{y},t)}\right)_{j=1}^{n}.$
@@ 9.1. Abstract log-concave interpolation on the real axis
Let $K_{j}(t):=\log L_{j}(t)$, $j=1,\ldots,n$,
(20) $J(t):=\begin{cases}\log w(t),\quad t\in\{x_{0},\ldots,x_{n}\},\\ -\infty,\quad t\in\mathbb{R}\setminus\{x_{0},\ldots,x_{n}\}.\end{cases}$
@@
We claim that the regularity set has the following form
(21) $R=\{\mathbf{y}\in\overline{S^{\infty}}:\ x_{j}<y_{j+1}<x_{j+1}\text{ for }j=0,\ldots,n-1\}.$
Indeed, let $y_{0}:=$-$\infty$, $y_{n}:=\infty$, and $j=0,\ldots,n$. By the singularity of the kernels, $m_{j}(\mathbf{y})>$-$\infty$ if and only if $J(t_{j})>$-$\infty$ for some point $t_{j}\in(y_{j},y_{j+1}).$ Thus, there are $n+1$ intervals $(y_{j},y_{j+1})$, and, by (20), there are exactly $n+1$ points $x_{j}$, where $J$ is finite. Consequently, $R$ is of the form (21).
@@ 9.2. Moving node Hermite–Fejér interpolation
For convenience, let $y_{0}:=$-$\infty$ and $y_{n+1}:=\infty$. Then, for any $\alpha_{0},\ldots,\alpha_{n}>0$ there are a unique $C>0$ and points $y_{1}<\ldots<y_{n}$ such that for the function
@@
The further proof is similar to the proof of Theorem 9.1 with $x_{j}$ replaced by $z_{j},\\ j=0,\ldots,n$. ∎